Isaac,Kamal and Arul have a total of $21.Isaac has $2 more than Kamal who has $5 more than Arul.How much does Arul have?
1
Expert's answer
2013-03-27T14:06:24-0400
Isaac, Kamal and Arul have a total of $21.Isaac has $2 more than Kamal who has $5 more than Arul.How much does Arul have?

Let Isaac, Kamal and Arul have I, K and A dollars each respectfully. Now let's make a system of equations and solve it:

I + K + A = 21 (1)
I - 2 = K (2)
K - 5 = A (3)

(3) ==> K = A + 5;

(2) ==> I = K + 2 = A + 7;

(1): A + 7 + A + 5 + A = 21 ==> 3A = 9 ==> A = 3;

K = 3 + 5 = 8;

I = 3 + 7 = 10.

So, Isaac has $10, Kamal has $8 and Arul has $3.